college algebra
use an equality and the five step process to solve the problem
translate
reword
carry-out
check
state
dg's plumbing and heating charges $50 plus $65 per hour for emergency service. Bill remembers being billed just over $450 for an emergency call. how long to the nearest hour was the plumber at bill's house?

Answer by mananth(16949)

dg's plumbing and heating charges $50 plus $65 per hour
billing $450 .
..
let the number of hours worked be $x
50+65x= 450
add -50 to both sides
50-50+65x=450-50
65x=400
divide by 65
65x/65 = 400/65
x= 6.15 hours
so plumber worked for about 6 hours
..
CHECK
50+6.15*65 = $449.75